# Energetics-based model for a diffusiophoretic motion of a deformable droplet

### v1

@@ Abstract
We construct a mathematical model for a diffusiophoretic motion of a deformable droplet, which is floating at a liquid surface and is driven by the surface tension gradient originating from the surface concentration field of the chemicals that are emitted from the droplet. We define the free energy of the system by including the surface and line energies. From the calculation of the functional of the free energy, we obtain a mathematical model for the diffusiophoretic motion with deformation. By only considering the deformation of the second mode, we explicitly derived the time-evolution equations for the translational motion and the elliptic deformation. We found three stable states: an immobile circular droplet, an immobile elliptically deformed droplet, and a moving droplet with the elliptic deformation in which the minor axis meets the motion direction, and discussed the transition between these three stable states.
@@ I Introduction
Under nonequilibrium conditions, particles and droplets can spontaneously move through the dissipation of chemical energy, which is often called self-propulsion. They have been intensively studied as a model for the motion of living organisms. They have also attracted interest as elements of active matter, in which spatio-temporal order at greater scales than those of elements emerges [1, 2, 3]. Many experimental systems on self-propulsion in physico-chemical systems have also been reported; oil droplets in surfactant aqueous solutions or aqueous droplets in oil with surfactant exhibit self-propulsion [4, 5, 6, 7, 8, 9, 10, 11]. Other examples are Quincke rollers, self-propelled plastic particles under an electric field [12, 13], and Janus particles, two-faced particles that move due to the temperature gradient or chemical potential gradient [14, 15, 16, 17]. Floating objects on liquid surfaces can exhibit self-propulsion due to the difference in surface tension, which are often called Marangoni surfers [18, 19, 20, 21, 22]. Droplet motions with adsorption and desorption of surfactants were discussed based on the low-Reynolds number hydrodynamics [23, 24, 25], and the self-propelled droplet on the solid surface in a surfactant aqueous solution is discussed based on the energetics [26, 27, 26, 28]. The solid Marangoni surfers have also been discussed based on hydrodynamics [29, 30] or the coupled system between Newtonian equations and reaction-diffusion equations [31, 32, 19, 33].
@@
In self-propulsion systems, the coupling between motion and shape has attracted interest. This is partly because the motion of living cells such as Keratocyte and Dictyostelium is strongly correlated with their shapes [34, 35, 36]. In also physico-chemical systems, the shape-dependent self-propulsion has been reported. Droplets that exhibit Marangoni surfing are often deformed and their shapes are strongly correlated with the direction of the motion [37, 38, 39, 40, 41, 42, 43, 44, 45]. This is because both self-propulsion and deformation have the same origin. For example, an oil droplet in a surfactant aqueous solution exhibits a self-propelled motion due to the Marangoni convection, which also causes ellipsoidal deformation [46]. A floating oil droplet at a water surface exhibits motion and deformation under a vertical vibration, both of which are led by Faraday waves [47]. Gel formation at the interface is also known to cause motion and deformation of the droplet [48] and a similar droplet behavior due to the actomyosin force generation was reported recently [49]. From a viewpoint of theoretical approach to the coupling between self-propulsion and deformation, there has also been a wide variety of mathematical modeling. One of the simplest models, so-called Ohta-Ohkuma model, described by coupled ordinary differential equations is proposed and analyzed based on the system symmetry [50, 51, 52]. This model only suggests the possible coupling manner and do not include the correspondence to the mechanism of the actual systems. On the other hand, some models have been developed based on insights from actual phenomena of the droplet motion and deformation [53, 54, 55]. Mathematical models that reproduce the motion and deformation of the cells have also been proposed [56, 57].
@@
We have investigated the motion of Marangoni surfers based on experiments and mathematical modeling comprising the Newtonian equations and reaction-diffusion equations. The Marangoni surfers are driven by surface tension gradient originating from the concentration profile of chemicals that are emitted from the surfers. They tend to escape from the chemicals and thus their motion mechanism is often called self-diffusiophoresis. We have considered such a self-diffusiophoretic particle with a given shape that emits surface-active chemicals and moves due to the surface tension gradient [58, 59, 60, 61, 33]. For example, theoretical analysis suggests that an elliptic Marangoni surfer should move in its minor axis direction, which is also confirmed by numerical simulations and experiments using a camphor particle. In these studies, the surfer has a rigid shape, and thus we do not need to discuss the change in shape. In contrast, deformable droplets can also move as Marangoni surfers. For example, a pure alcohol droplet like pentanol and hexanol on an alcohol aqueous solution is known as the examples of such deformable moving droplets [37, 38, 62]. These experimental results clearly demonstrate the coupling between motion and deformation. Some of the authors proposed a mathematical model based on the energy minimization but the cost for numerical simulation was quite high [54]. Therefore, a mathematical model that well reproduces experimental results and enables us to discuss the mechanism from a viewpoint of system symmetry is awaited. Recently, some of the authors proposed a mathematical model adopting a phase-field description for the droplet shape [55]. The model succeeded in reproducing self-propulsion coupled with deformation but the relation to the simple models like the Ohta-Ohkuma model [50, 51], is still unveiled. Through the discussion of the correspondence between these models, we believe that we can attain the essential mechanism of the diffusiophoretic motion with deformation.
@@
In the present study, thus, we propose a mathematical model for the diffusiophoretic deformable droplet, in which a droplet shape is described by the Fourier series expansion. Our model corresponds to a system with a floating deformable droplet that releases surface-active chemicals and generates their concentration field. The motion and deformation of the droplet are driven by the surface tension gradient originating from the concentration field of the chemicals. The system dynamics are partly derived using a variational principle of the free energy of the system. In the manuscript, we first show our mathematical model together with brief derivation of it in Sect. II. Then, we limit the case in which a droplet shows translational motion and the 2-mode deformation. We show the explicit form of the model and perform numerical simulations in Sect. III. Further, we perform theoretical analysis using the perturbation method and compare with the simulation results in Sect. IV. We discuss the results by numerical simulation and theoretical analyses in Sect. VI and summarize our work in Sect. VII. Since the present model is derived naturally from the free energy of the system, the model can be a universal one for the diffusiophoretic motion of a deformable object in a two-dimensional space.
@@ II Modeling
Here, $D$ is an effective diffusion coefficient that includes the hydrodynamic effect [63, 64, 65], $\alpha$ is the evaporation and dissolution rate, and $S(\bm{r};\left\{a_{k}\right\},\left\{b_{k}\right\})$ are the supply rate of the surface-active chemicals from the droplet. The term $S$ is explicitly expressed as
@@
Here, $S_{0}$ is the supply rate of the surface active chemicals from the droplet, and $A$ is the base area of the droplet, which is explicitly obtained as $A=\pi R^{2}+\mathcal{O}(\epsilon^{2})$.
@@ III Numerical simulation in the case with 2-mode deformation
Based on the model derived above, we performed a numerical simulation. For the correspondence to the theoretical analyses described in the following section, we only consider the 2-mode deformation. In order to minimize the effect of discretization, we define a smooth function representing the area inside the droplet as
@@
For the numerical simulation, we adopted an explicit method for the time development with spatial mesh $\Delta x=0.025$ and time step $\Delta t=0.0001$. The system size was set to be $7\times 7$ with the periodic boundary condition. The smoothing parameter was set to be $\sigma=0.03$. We checked that the results did not change significantly if we slightly changed the value of $\sigma$. We fixed $R=1$ and $\eta_{2}=0.1$, and varied $\eta_{t}$ and $\kappa_{2}$ as the parameters. The initial conditions for $u$ and $\bm{r}_{c}$ were fixed as $u=0$ and $\bm{r}_{c}=\bm{0}$, and those for $d\bm{r}_{c}/dt$, $a_{2}$, and $b_{2}$ were varied.
@@
Figure 1: Snapshots at $t=1000$ (left panels) and time series (right panels) of the speed $v=\left|d\bm{r}_{c}/dt\right|$ (red) and the deformation magnitude $s=\sqrt{{a_{2}}^{2}+{b_{2}}^{2}}$ (blue) for the three typical cases. (a) Immobile circular (IC) droplet at $\kappa_{2}=0.12$ and $\eta_{t}=0.085$. (b) Immobile deformed (ID) droplet at $\kappa_{2}=0.116$ and $\eta_{t}=0.085$. (c) Mobile deformed (MD) droplet at $\kappa_{2}=0.112$ and $\eta_{t}=0.085$. The droplet moves downward as the orange arrow indicates. Grayscale tone displays the concentration field $u$. The initial condition was set as $d\bm{r}_{c}/dt=\bm{e}_{x}+0.1\bm{e}_{y}$, $a_{2}=0.1$, $b_{2}=0.01$.
@@
In Fig. 1, we show a typical example of the droplet dynamics. For great $\eta_{t}$ and $\kappa_{2}$, we observed an immobile circular (IC) droplet, which kept a circular shape and did not move as shown in Fig. 1(a). For smaller $\kappa_{2}$ and greater $\eta_{t}$, we observed an immobile deformed (ID) droplet, which deformed into an elliptic or peanut shape but did not move, as shown in Fig. 1(b). For smaller $\eta_{t}$, we observed a moving deformed (MD) droplet, which moved in a certain direction with an elliptic deformation whose direction of the minor axis met the direction of motion as shown in Fig. 1(c).
Figure 2: Phase diagram on the $\eta_{t}$-$\kappa_{2}$ plane. The other parameters were set to be $\eta_{2}=0.1$ and $R=1$. The red, green, and blue points correspond to an immobile circular (IC) droplet, an immobile deformed (ID) droplet, and a mobile deformed (MD) droplet. The cyan points show the bistability between IC and MD droplets. The orange points show the other cases, in which we could classify into none of the above.
In Fig. 2, we show a phase diagram of stable states on the $\eta_{t}$-$\kappa_{2}$ plane. The other parameters were fixed at $\eta_{2}=0.1$ and $R=1$. We calculated with the following four initial conditions (i) $d\bm{r}_{c}/dt=\bm{e}_{x}+0.1\bm{e}_{y}$, $a_{2}=0.1$, $b_{2}=0.01$, (ii) $d\bm{r}_{c}/dt=\bm{e}_{x}+0.01\bm{e}_{y}$, $a_{2}=-0.25$, $b_{2}=0.002$, (iii) $d\bm{r}_{c}/dt=0.01\bm{e}_{x}+0.001\bm{e}_{y}$, $a_{2}=0.1$, $b_{2}=0.01$, and (iv) $d\bm{r}_{c}/dt=0.01\bm{e}_{x}+0.001\bm{e}_{y}$, $a_{2}=0.01$, $b_{2}=0.001$, where the initial conditions (i) and (ii) have a large velocity and deformation with different direction relations, (iii) has a small velocity and a large deformation, and (iv) has a small velocity and deformation. In the phase diagram, the IC droplet was stable for $\eta_{t}\gtrsim 0.077$ and $\kappa_{2}\gtrsim 0.119$. With a decrease in $\kappa_{2}$, the IC droplet became unstable and an ID droplet became stable. In contrast, with a decrease in $\eta_{t}$, an IC droplet became unstable and the MD droplet was realized. Close to the transition points between IC and MD droplets, we observed the bistable state of them. The orange points around $\eta_{t}\simeq 0.077$ and $\kappa_{2}\gtrsim 0.153$ had a finite speed with a small deformation. They could be classified into another type like “mobile circular droplets”, but since we only observed them at few parameter values and found they showed slight deformation, we showed them as “cannot be classified”.
In Fig. 3, the one-dimensional bifurcation diagrams are shown. Here, we adopted symmetric initial conditions (i’) $d\bm{r}_{c}/dt=\bm{e}_{x}$, $a_{2}=-0.1$, $b_{2}=0.0$, and (ii’) $d\bm{r}_{c}/dt=0.001\bm{e}_{x}$, $a_{2}=-0.1$, $b_{2}=0$ to clearly show the bifurcation structures of the solutions. In Fig. 3(a), the bifurcation diagram with respect to $\kappa_{2}$ at $\eta_{t}=0.085$ is shown. The deformation magnitude $s$ rose at $\kappa_{2}\lesssim 0.118$. For $0.115\lesssim\kappa_{2}\lesssim 0.118$, the droplet speed was zero, which indicated an ID droplet. For $\kappa_{2}\lesssim 0.115$, another branch with finite speed and deformation amplitude discontinuously appeared, which indicated an MD droplet. The transition from an IC droplet to an ID droplet at $\kappa_{2}\simeq 0.118$ seems to be a supercritical pitchfork bifurcation. In contrast, the transition from an ID droplet to an MD droplet seems discontinuous since there exists a bistable state. In Fig. 3(b), the bifurcation diagram with respect to $\eta_{t}$ for $\kappa_{2}=0.13$ is shown. The droplet speed $v$ and the deformation amplitude simultaneously rose at $\eta_{t}\lesssim 0.079$. The IC droplet became unstable at $\eta_{t}\simeq 0.076$. The transition from an IC droplet to an MD droplet seemed discontinuous, which implies a subcritical pitchfork bifurcation.
Figure 3: Phase diagrams with constant $\eta_{t}$ and $\kappa_{2}$, which correspond to the vertical and horizontal lines, respectively. (a) Phase diagram at $\eta_{t}=0.085$. Speed $v$ and deformation $s$ are plotted depending on $\kappa_{2}$. (b) Phase diagram at $\kappa_{2}=0.13$ depending on $\eta_{t}$. Speed $v$ and deformation $s$ are plotted depending on $\kappa_{2}$. The red, blue, and green points represent immobile circular (IC), immobile deformed (ID), and mobile deformed (MD) droplets. The vertical dotted lines are drawn to indicate the transition points and correspondence between $v$ and $s$.
@@
For even smaller $\eta_{t}$ or $\kappa_{2}$, the calculation diverged due to $s$. This is because our modeling is based on the assumption that the deformation is sufficiently small as the first-order perturbation from a circle. Actually, $s$ is well-defined in the range of $\left|s\right|<1$.

### v2

@@ Abstract
We construct a mathematical model for a diffusiophoretic motion of a deformable droplet, which is floating at a liquid surface and is driven by the surface tension gradient originating from the surface concentration field of the chemicals that are emitted from the droplet. We define the free energy of the system by including the surface and line energies. From the calculation of the functional of the free energy, we obtain a mathematical model for the diffusiophoretic motion with deformation. By only considering the deformation of the second mode, we explicitly derive the time-evolution equations for the translational motion and the elliptic deformation. There are three stable states: an immobile circular droplet, an immobile elliptically deformed droplet, and a moving droplet with the elliptic deformation in which the minor axis meets the motion direction, and we discuss the transition between these three stable states.
@@ I Introduction
Under nonequilibrium conditions, particles and droplets can spontaneously move through the dissipation of chemical energy, which is often called self-propulsion. They have been intensively studied as a model for the motion of living organisms. They have also attracted interest as elements of active matter, in which spatio-temporal order at greater scales than those of elements emerges [1, 2, 3]. Many experimental systems on self-propulsion in physico-chemical systems have also been reported; crawling droplets on a solid substrate [4, 5, 6, 7, 8], oil droplets in surfactant aqueous solutions or aqueous droplets in oil with surfactant exhibit self-propulsion [9, 10, 11, 12, 13, 14, 15, 16]. Other examples are Quincke rollers, self-propelled plastic particles under an electric field [17, 18], and Janus particles, two-faced particles that move due to the temperature gradient or chemical potential gradient [19, 20, 21, 22]. Floating objects on liquid surfaces can exhibit self-propulsion due to the difference in surface tension, which are often called Marangoni surfers [23, 24, 25, 26, 27]. Droplet motions with adsorption and desorption of surfactants were discussed based on the low-Reynolds number hydrodynamics [28, 29, 30], and the self-propelled droplet on the solid surface in a surfactant aqueous solution is discussed based on the energetics [31, 32, 31, 33]. The solid Marangoni surfers have also been discussed based on hydrodynamics [34, 35] or the coupled system between Newtonian equations and reaction-diffusion equations [36, 37, 24, 38].
@@
In self-propulsion systems, the coupling between motion and shape has attracted interest. This is partly because the motion of living cells such as Keratocyte and Dictyostelium is strongly correlated with their shapes [39, 40, 41]. In physico-chemical systems, the shape-dependent self-propulsion has also been reported. Droplets that exhibit Marangoni surfing are often deformed and their shapes are strongly correlated with the direction of the motion [42, 43, 44, 45, 46, 47, 48, 49, 50]. This is because both self-propulsion and deformation have the same origin. For example, an oil droplet in a surfactant aqueous solution exhibits a self-propelled motion due to the Marangoni convection, which also causes ellipsoidal deformation [51]. Gel formation at the interface is also known to cause motion and deformation of the droplet [52] and a similar droplet behavior due to the actomyosin force generation was reported recently [53]. From a viewpoint of theoretical approach to the coupling between self-propulsion and deformation, there has also been a wide variety of mathematical modeling. One of the simplest models, so-called Ohta-Ohkuma model, described by coupled ordinary differential equations is proposed and analyzed based on the system symmetry [54, 55, 56]. This model only suggests the possible coupling manner and do not include the correspondence to the mechanism of the actual systems. On the other hand, some models have been developed based on insights from actual phenomena of the droplet motion and deformation [57, 58, 59]. Mathematical models that reproduce the motion and deformation of the cells have also been proposed using the phase-field model [60, 61, 62] and the Fourier series expansion [63].
@@
We have investigated the motion of Marangoni surfers based on experiments and mathematical modeling comprising the Newtonian equations and reaction-diffusion equations. The Marangoni surfers are driven by surface tension gradient originating from the concentration profile of chemicals that are emitted from the surfers. The surfers tend to escape from the chemicals and thus their motion mechanism is often called self-diffusiophoresis. We have considered such a self-diffusiophoretic particle with a given shape that emits surface-active chemicals and moves due to the surface tension gradient [64, 65, 66, 67, 38]. For example, theoretical analysis suggests that an elliptic Marangoni surfer should move in its minor axis direction, which is also confirmed by numerical simulations and experiments using a camphor particle. In these studies, the surfer has a rigid shape, and thus we do not need to discuss the change in shape. In contrast, deformable droplets can also move as Marangoni surfers. For example, a pure alcohol droplet like pentanol and hexanol on an alcohol aqueous solution is known as the examples of such deformable moving droplets [42, 43, 68]. These experimental results clearly demonstrate the coupling between motion and deformation. Some of the authors proposed a mathematical model based on the energy minimization but the cost for numerical simulation was quite high [58]. Therefore, a mathematical model that well reproduces experimental results and enables us to discuss the mechanism from a viewpoint of system symmetry is awaited. Recently, some of the authors proposed a mathematical model adopting a phase-field description for the droplet shape [59]. The model succeeded in reproducing self-propulsion coupled with deformation but the relation to the simple models like the Ohta-Ohkuma model [54, 55], is still to be unveiled. Through the discussion of the correspondence between these models, we believe that we can attain the essential mechanism of the diffusiophoretic motion with deformation.
@@
In the present study, thus, we propose a mathematical model for the diffusiophoretic deformable droplet, in which a droplet shape is described by the Fourier series expansion. Our model corresponds to a system with a floating deformable droplet that releases surface-active chemicals and generates their concentration field. The motion and deformation of the droplet are driven by the surface tension gradient originating from the concentration field of the chemicals. The system dynamics are partly derived using a variational principle of the free energy of the system. In the manuscript, we first show our mathematical model together with brief derivation of it in Sect. II. Then, we limit the case in which a droplet shows translational motion and the 2-mode deformation. We show the explicit form of the model and show the numerical simulation results in Sect. III. Further, we adopt the perturbation method and obtain the reduced model in Sect. IV. Then, we perform a linear stability analysis of the obtained reduced model and compare with numerical simulation results in Sect. V Finally, we summarize our work and show the future direction in Sect. VI. Since the present model is derived naturally from the free energy of the system, the model can be a universal one for the diffusiophoretic motion of a deformable object in a two-dimensional space.
@@ II Modeling
Here, $\alpha$ is the evaporation and dissolution rate. The transport of chemicals should be induced by the Marangoni effect [69] and it should be described by a nonlinear term. However, it is shown that the spreading effect of the concentration field by the Marangoni convection can be approximated using a linear diffusion term $D\nabla^{2}u$ with an effective diffusion coefficient $D$ if the system is close to the bifurcation point from a steady state [70, 71, 72]. Thus, we adopt the effective linear diffusion term. $S(\bm{r};\left\{a_{k}\right\},\left\{b_{k}\right\})$ are the supply rate of the surface-active chemicals from the droplet. The term $S$ is explicitly expressed as
@@
Here, $S_{0}$ is the supply rate of the surface active chemicals from the droplet, and $A$ is the base area of the droplet, which is explicitly obtained as $A=\pi R^{2}+\mathcal{O}(\epsilon^{2})$. In our model, we assume that the chemicals are released not only from the droplet periphery but also from the whole droplet region. If we only introduce the supply of chemicals at the periphery, we need to consider the reaction-diffusion equation that is valid only for the outside of the droplet region. In such a case, we need to include a Neumann boundary condition at the moving droplet periphery. By introducing the approximation that the chemicals are supplied in the whole droplet region, we obtain a time evolution equation that is valid for the whole two-dimensional space and can approximate the concentration dynamics around the droplet.
@@ III Numerical simulation in the case with 2-mode deformation
Based on the partial differential equation (PDE) model derived above, we perform a numerical simulation. For the correspondence to the theoretical analyses described in the following section, we only consider the 2-mode deformation. In order to minimize the effect of discretization, we define a smooth function representing the area inside the droplet as
@@
For the numerical simulation, we adopt an explicit method for the time development with spatial mesh $\Delta x=0.025$ and time step $\Delta t=0.0001$. The system size is set to be $7\times 7$ with the periodic boundary condition. The smoothing parameter is set to be $\sigma=0.03$. It is confirmed that the results do not change significantly if we slightly change the value of $\sigma$. We fix $R=1$ and $\eta_{2}=0.1$, and vary $\eta_{t}$ and $\kappa_{2}$ as the parameters. The initial conditions for $u$ and $\bm{r}_{c}$ are fixed as $u=0$ and $\bm{r}_{c}=\bm{0}$, and those for $d\bm{r}_{c}/dt$, $a_{2}$, and $b_{2}$ are varied.
@@
Figure 1: Numerical simulation results obtained from the PDE model in Eqs. (35) to (43). Snapshots at $t=1000$ (left panels) and time series (right panels) of the speed $v=\left|d\bm{r}_{c}/dt\right|$ (red) and the deformation magnitude $s=\sqrt{{a_{2}}^{2}+{b_{2}}^{2}}$ (blue) are shown for the three typical cases. (a) Immobile circular (IC) droplet at $\kappa_{2}=0.12$ and $\eta_{t}=0.085$. (b) Immobile deformed (ID) droplet at $\kappa_{2}=0.116$ and $\eta_{t}=0.085$. (c) Mobile deformed (MD) droplet at $\kappa_{2}=0.112$ and $\eta_{t}=0.085$. The droplet moves downward as the orange arrow indicates. Grayscale tone displays the concentration field $u$. The initial condition is set as $d\bm{r}_{c}/dt=\bm{e}_{x}+0.1\bm{e}_{y}$, $a_{2}=0.1$, $b_{2}=0.01$.
@@
In Fig. 1, we show a typical example of the droplet dynamics. For great $\eta_{t}$ and $\kappa_{2}$, we observe an immobile circular (IC) droplet, which keeps a circular shape and does not move as shown in Fig. 1(a). For smaller $\kappa_{2}$ and greater $\eta_{t}$, we observe an immobile deformed (ID) droplet, which deforms into an elliptic or peanut shape but does not move, as shown in Fig. 1(b). For smaller $\eta_{t}$, we observe a moving deformed (MD) droplet, which moves in a certain direction with an elliptic deformation whose direction of the minor axis meets the direction of motion as shown in Fig. 1(c).
Figure 2: Numerical simulation results obtained from the PDE model in Eqs. (35) to (43). (a) Two-dimensional phase diagram on the $\eta_{t}$-$\kappa_{2}$ plane. The other parameters are set to be $\eta_{2}=0.1$ and $R=1$. The red, green, and dark blue points correspond to an immobile circular (IC) droplet, an immobile deformed (ID) droplet, and a mobile deformed (MD) droplet. The cyan points show the bistability between IC and MD droplets, the yellow points show that between IC and ID droplets, and the gray points show that between IC and mobile circular (MC) droplet. The brown points represent the case that we can classify into none of the above (Three-or-more final steady states due to the slow convergence of the system). (b) One-dimensional bifurcation diagrams with constant $\eta_{t}$ or $\kappa_{2}$, which are indicated as arrows at the top and right sides of the plot in panel (a). Speed $v$ and deformation $s$ are plotted. (i,ii) $\kappa_{2}$ is varied at (i) $\eta_{t}=0.082$ and (ii) $\eta_{t}=0.090$. (iii-vi) $\kappa_{2}$ is varied at (iii) $\kappa_{2}=0.160$, (iv) $\kappa_{2}=0.130$, (v) $\kappa_{2}=0.117$, and (vi) $\kappa_{2}=0.100$. The red, green, and dark blue lines represent immobile circular (IC), immobile deformed (ID), and mobile deformed (MD) droplets. For the MD droplet, it moves in its minor-axis direction. The vertical dotted lines are drawn to indicate the transition points and correspondence between $v$ and $s$.
In Fig. 2(a), we show a two-dimensional phase diagram of stable states on the $\eta_{t}$-$\kappa_{2}$ plane. The other parameters are fixed at $\eta_{2}=0.1$ and $R=1$. We start with the following four initial conditions (i) $d\bm{r}_{c}/dt=\bm{e}_{x}+0.1\bm{e}_{y}$, $a_{2}=0.1$, $b_{2}=0.01$, (ii) $d\bm{r}_{c}/dt=\bm{e}_{x}+0.01\bm{e}_{y}$, $a_{2}=-0.25$, $b_{2}=0.002$, (iii) $d\bm{r}_{c}/dt=0.01\bm{e}_{x}+0.001\bm{e}_{y}$, $a_{2}=0.1$, $b_{2}=0.01$, and (iv) $d\bm{r}_{c}/dt=0.01\bm{e}_{x}+0.001\bm{e}_{y}$, $a_{2}=0.01$, $b_{2}=0.001$, where the initial conditions (i) and (ii) have a large velocity and deformation with different direction relations, (iii) has a small velocity and a large deformation, and (iv) has a small velocity and deformation. From each initial condition, the numerical simulation is run until it reaches a steady state, where we consider the system reaches a steady state if maximum time difference of $v_{x}$, $v_{y}$, $a_{2}$, and $b_{2}$ are less than $10^{-6}$ per 0.1 time unit. We classify the final steady-state using $v=\sqrt{{v_{x}}^{2}+{v_{y}}^{2}}$ and $s=\sqrt{{a_{2}}^{2}+{b_{2}}^{2}}$ as follows: an IC droplet in case that $v\leq 0.01$ and $s\leq 0.01$, an ID droplet in case that $v\leq 0.01$ and $s>0.01$, MD droplet in case that $v>0.01$ and $s>0.01$, and MC (mobile circular) droplet in case that $v>0.01$ and $s\leq 0.01$. If the final steady states are the same from all four initial conditions, then we consider the system to have a single final state. If there are two final steady states depending on the choice of the initial conditions, then the system is considered to be in a bistable state. If there are three-or-more final steady states, then the system is “not classified”.
In the phase diagram, the IC droplet is stable for $\eta_{t}\gtrsim 0.077$ and $\kappa_{2}\gtrsim 0.119$. With a decrease in $\kappa_{2}$, the IC droplet becomes unstable and an ID droplet becomes stable. In contrast, with a decrease in $\eta_{t}$, an IC droplet becomes unstable and the MD droplet is realized. Close to the transition points between IC and MD droplets and those between ID and MD droplets, we observe the bistable states of them, respectively.
The gray points around $\eta_{t}\simeq 0.077$ and $\kappa_{2}\gtrsim 0.153$ correspond to the bistable state between the IC droplet and the mobile circular (MC) droplet, where the droplet has a finite speed with a small deformation. This steady state is classified into the MC droplet since the amplitude of deformation is quite small. Thus, we consider that the states corresponding to the gray points are the intermediate region between IC and MD droplets.
In Fig. 2(b), the one-dimensional bifurcation diagrams are shown. Here, we adopt symmetric initial conditions (i’) $d\bm{r}_{c}/dt=\bm{e}_{x}$, $a_{2}=-0.1$, $b_{2}=0.0$, and (ii’) $d\bm{r}_{c}/dt=0.001\bm{e}_{x}$, $a_{2}=-0.1$, $b_{2}=0$ to clearly show the bifurcation structures of the solutions. In Fig. 2(b)-(i,ii), the bifurcation diagrams with respect to $\kappa_{2}$ at $\eta_{t}=0.082$(i) and 0.0900.090(ii) are shown. For $\eta_{t}=0.082$, the bistability of the MD and ID droplets and that of the MD and IC droplets are shown. The transition between the IC and ID droplets is continuous. For $\eta_{t}=0.090$, the IC droplet continuously changes to the ID droplet and then to the MD droplet with a decrease in $\kappa_{2}$. In Fig. 2(b)-(iii-vi), the bifurcation diagrams with respect to $\eta_{t}$ for $\kappa_{2}=0.160$(iii), 0.1300.130(iv), 0,1170,117(v) and 0.1000.100(vi) are shown. The IC droplet continuously changes to the MD droplet with a decrease in $\eta_{t}$ for $\kappa_{2}=0.160$, while the transition between the IC and MD droplets is discontinuous and has a hysteresis. Also for the transition between the ID and MD droplets, discontinuous and continuous transitions are seen for $\kappa_{2}=0.117$ and $0.100$, respectively.
@@
For even smaller $\eta_{t}$ or $\kappa_{2}$, the calculation can diverge due to $s$. This is because our modeling is based on the assumption that the deformation is sufficiently small as the first-order perturbation from a circle. Actually, $s$ is well-defined in the range of $\left|s\right|<1$.